FEL Upgrade Project Weekly Brief
June 25-29, 2001

This week we completed a second full week of experimental operations for the parametric study on laser-materials damage (i.e.,"lethality") funded by the Joint Technology Office.  See the "Operations" section for details.

On June 26 we hosted a review of the FEL program for SURA's Maritime Technical Advisory Committee (MTAC), chaired by VADM (ret.) Al Baciocco.

Note: Next week most of the FEL team will be taking a well earned summer break after a busy half year of Upgrade design activities and FEL operations.  Except for some orbit studies planned by our intrepid beam transport designer (Dave Douglas), there will be no operations next week.  Full operations and the next weekly report will reappear the week of July 9th.

We used the occasion of the MTAC meeting this week to present a first order cut on our conceptual design for a
100 kW upgrade. Also the committee was a useful sounding board for the preliminary findings from the Navy high power FEL workshop held earlier ion the month, which were presented by the workshop chairman Alan Todd.

The project monthly report for May was completed and distributed to the DOE and ONR contracting officers.  As of the end of May the project is ahead of schedule (49% complete vs 44% scheduled).  We are on track for having the first phase of the project fully obligated by September.  A negative cost variance of ~$1M is presently showing but this is largely an artifact of having to obligate funds early in the project in order to ensure full obligation and costing by the project end date.

WBS 3 (Beam Physics):
WBS 3 is off the air due to CEBAF program deputy duties.  All beam physics requests will be processed, in the order received, starting July 23. If you have a beam physics emergency, please contact F. Dylla, x 7450.

WBS 4 (Injector):
The leak in the mid-voltage kovar ring was fixed.  All of the components for the ceramic stack  are going through final cleaning before the stack is welded together.  Continued refining the gun chamber drawings which should be complete with the exception of the screens and associated mounts which are 50% complete.  The materials for the gun chamber are in house.

Gun HVPS - The Gun HVPS has been received and is in the vault.  The design of the new HVPS Tank and HV Transmission Line has started with a new mechanical engineer.

WBS 6 (RF):
Zone 3 - B. Nichols was assigned to complete this zone by the end of August.  The Circulators are staged for installation.  The HVPS is due to be received in 2 - 3 weeks from Hipotronics.

Zone 4 - No progress on completing the punch list.  New RF Control Modules are being built, but none have been received to date.

Injector Klystrons - CPI believes they have resolved the helium leak test questions are proceeding to complete the three 100 kW klystrons.  No date has been set for factory tests.

Quarter HVPS's - Quotes are due today for the transformers and inductors to upgrade the HVPS's.  Receiving information for the rectifier assemblies and SCR controllers.

WBS 8 (Instrumentation):
Provided user support while running.  Also purchased an audio system to allow audio recording of experiments as well as video.

New VME Timing/Delay boards are on order.

Mike and Patrick have been working on the computer in User Lab 4 to get it connected to the user LAN.  Mike has added a search feature to the Task System on the laser website.  The next step in this process will be the ability to search the FLOG in a read only format.  Patrick has programmed the rabbit board to operate servo motors remotely and built a chassis for it.  The next step will be to access the LAN for control from off site.

The pulse truncater for the Drive Laser has been connected.

James and Dan have completed the video patch panel for the hutch in User Lab 3.

The spare VME Interface modules are complete.  The Isolation modules are waiting on parts.

OBPM work continues.  A tracking system using serial numbers has been established for better control of testing of the detectors.  The absolute value/sample and hold board drawing is in EECAD.  Prototype boards should be out soon.

Drawings in EECAD: Molectron Quad Buffer Amplifier, Fabrication, Schematic and Assembly; Analog/Differential Driver Board Schematic; Pyro-electric Detector Buffer Amplifier, Schematic, Fabrication and Assembly; 32 Channel Sample/Hold Board (VME) Schematic, Assembly and Fabrication; Ion Pump Power Supply Controller Board Schematic; Beam Viewer Chassis Test Box Schematic, Assembly and Fabrication; LSS PLC Test Box Schematic, Assembly and Fabrication; User Shutter Control Box Wiring Diagram; MPS System Diagram; Absolute Value/Sample and Hold Board Schematic, Assembly and Fabrication; LSS Wiring Diagram- Revision.

WBS 9 (Transport):
Injector Dipoles (DU/DV)
o  The DULY is working on the revisions to the drawings of the Injector Dipoles (GU & GV).
Arc Dipoles (GY, GX, GQ)
o  Advanced Energy Systems finished the details of the leads, bus bars, manifold, alignment cartridge mounts and
    lead cover of the Arc Bend Dipole (GX) and have submitted all drawings of this magnet for checking.
o  In magnetic modeling pursuit of flat longitudinal field, the use of micrometric decreases in gap size in bands near
    the edges succeeded and we will be going on to now establish the effective length by adjusting the position of the
    end faces.  This will enable us to then fix the gap on the 180 Degree Dipole (GY).
o  AES started design of the last magnet of the series, the Reverse Bend (GQ).  It will be able to feed directly off
    the design of the Regular Bend (GX), so rapid progress is expected.
QX (3 inch quad)
o  The core manufacturing contract was placed.
QT (Trim Quad)
o  Final magnetic modeling reduced the width by 2 inches.  A last pass at the drawings incorporated the changes.
o  Robin Wines continued work on the 3D model in TOSCA.  Drafting design started.
o  Dave Douglas reduced the field required of the Octupole considerably.  A task order for DULY Research to
    start magnetic modeling and design was composed.
o  The Dipole Steel order is still awaiting vendor response to a revised flatness tolerance that is due today.
Beam Line and Vacuum
o  Substantial work was done on the Interface Control Document for the task definition for the Arc and Optical
    Chicane chambers.
o  The copper order for all the major dipoles was placed last week.
o  The layout drawings for the girders between cryomodules and for the return leg were completed and submitted
    to all interested parties for review.

WBS 10 (Wiggler):
Analysis of the dispersion section measurements with a single turn correction coil verify that the field integral can be zeroed to +/- 10 G-cm over the full range of the magnet strength.  An alternate scheme will be tested as well.  A scheme for adding a correction coil to the dispersion section was developed.  The correction coils for the wigglers are done and will be mounted next week.

A test weld piece was made of a viewer cross for the wiggler chamber.  Despite using a fusion weld technique, the measured permeability was approximately 1.5 at the weld bead.  Since this is much higher than for the IR Demo we decided to measure the effect of the weld on the dispersion section field.  Measurements were made which indicated changes to the field of about 1 G near the weld. Since the electron beam does not get very near the weld and since a change of 1 G over a few centimeters would be acceptable anyway, we decided that the weld technique was acceptable.  The full chambers will now be fabricated.  Most of the rest of the pieces for the viewer camera assemblies and the vacuum chamber piece parts are now fabricated and in house.

WBS 11 (Optics):
Again, most of the optics physics team (Michelle and Joe) was devoted to the execution of the JTO experiments this week. With regard to the Upgrade:
    -detailing continues on the internal mounting hardware in the optical cavity chambers
    -we began design of the insertable mirror for the diagnostics mounted between the outcoupler chamber and the
      collimator.  In the process, we began exploring a suggestion that we use two insertable viewers (or ultraviewers),
      a design which will result in cost savings.

Work proceeded at a rapid pace taking ablation data for the JTO program. As of this writing we were completing data on Suite 40 of 6 test samples each. A large number of different sample types have been tested and FEL pulsed and CW operation at power levels to 1 kW and above. Substantial amounts of data have been acquired and will require some significant time to fully reduce and analyze.

On one day we ran for 1 kW for over 6 hours. We have also seen FEL efficiency run as high as >2% due to good optimization of the beam transport. This bodes well for our Upgrade plans which only require 0.8%.  Unfortunately with such a high CW power running uninterrupted the un-cooled mirrors in the optical transport line eventually heated to unacceptable levels and oxidized a reflective coating on one. Replacement of that mirror brought us back to full power operation the next morning.  As a consequence of that experience we are limiting our high power CW running to periods of an hour or two until we can attach some temperature stabilization clamps to the most stressed components.  The machine ran well during all this activity and we are able to produce 3 more runs of nanotube production as well as parasitically map the Thompson X-ray production while the JTO data was being acquired. The fluxes produced there were also high enough to saturate that detector under many conditions.  We are adding tighter collimation.